

********************************************************************
UARS WEEKLY STATUS REPORT - 07 April 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
On UARS day 93, SUSIM was originally scheduled to be off, due to
ongoing spacecraft power constraints. However, a change in the
schedule made it possible for SUSIM to remain powered. Unfortunately,
this came too late for SUSIM to submit modified plans to allow solar
viewing for UARS day 93.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. At the
present time, PEM energy deposition is not being generated. Recent data
configuration changes have affected PEM software causing multiple failures.
These failures are being reviewed by the PEM team. In order to correct for new
telemetry requirements, PEM will need to update processing code with a patch
delivery. Updated HEPS calibration parameters were finalized and are ready for
delivery. Updated MEPS quality calibration files are being generated which are
used during Level 3 product generation. Level 3 software errors, discovered by
Dr. Dean Pesnell, have been fixed; his suggestions have been incorporated with
in the PEM production code. Both updated HEPS and MEPS calibration data, along
with the Level 3 code changes, will influence the in-situ energy deposition
predictions, so new versions of PEM energy deposition Level 3 files will be
generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating region. Science data results continue to be
excellent.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the Monthly Status Review
Meeting, held on 5 April.
A revised instrument operations schedule was prepared
and published. With the current state of the
atmosphere, the MLS team offered to cancel the MLS
observations scheduled for 6 through 10 April.
scheduled MLS operating time. ACRIM, SOLSTICE,
SUSIM and WINDII are now operating through this
period.
Routine processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
STATUS:
- The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day HAL (4.82) PROD (2000) (1R-1C=0L)
- The CDHF will load February 2000 current day DAAC data in April 2000.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, SUSIM,
Solstice, and WINDII.
- Testing HRDI V4.4 on the Rehost System continues.
- Preparing to test MLS V4.22 on the Rehost System.
- Level0 Ingest, Quicklook Ingest, and NRT are being tested.
- Work instructions continue to be developed and revised. OPS personnel
are currently reviewing all work instructions for comments/corrections.
***********
FOT
***********
REPORTING PERIOD
This report covers 25 March, 2000 (Orbit 46654, GMT Day 085) through 31 March
2000 (Orbit 46758, GMT Day 091).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 086/0113 46669 On
PEM 087/0115 46684 AXIS On
PEM 087/0318 46685 AXIS-2 Off
MLS 090/2344 46743 Off
HALOE 091/0024 46763 On
The beta angle ranged between the angles of 13 degrees to 35 degrees for this
report period. The beta angle is now increasing toward a maximum angle of
+49.0 degrees on 08 April 2000 (DOY 099).
Spacecraft battery 1, 2, and 3 performance data is not available for this
report period.
The clock error ranged from 12.3 to -12.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 07
% HRDI = 05
% PEM = 02
% MLS = 02
% HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 00
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
7 Transponder A: 00/021 - 2287.496849
7 Transponder B: 00/060 - 2287.499207
GROUND SYSTEM OPERATION
Zero (0) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/090 show the observatory data loss to be 301 hours, 12
minutes, 33.513 seconds (an increase of 29 minutes 41.480 seconds since the
last report period).
The increase of 29 minutes 41.480 seconds occurred over five (5) days:
7 14 minutes 05.144 seconds on Day 086. Reason unknown.
7 05 minutes 07.976 seconds on Day 087. Data hits.
7 02 minutes 03.240 seconds on Day 088. Data hits.
7 03 minutes 00.920 seconds on Day 089. Data hits.
7 05 minutes 24.200 seconds on Day 090. Data hits.
These calculations have been recalculated to include 1999/346 through 2000/006
statistics using scheduled times.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Simulator support programmers worked on getting the UARS Simulator up to
date. The FOT delivered the TMONUs, RTS and Ephemeris needed for the
generation of the new test scenario. They encountered some problems but
believe that it will be ready to test the new TMONUs by next week.
The Relative Time Sequence mentioned above was generated in the CMS without
error. It is ready for uplink following the TMON testing and approval.
MEETINGS
The FOT did not attend any meetings this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for April 25, 2000 (DOY 116). This
maneuver will be a Reverse to Forward Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 31 March 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46656 085/0518 00-118 NBTR NBTR OPS
46669 086/0113 00-119 MLS POWER ON
46681 086/2154 00-120 NBTR NBTR OPS
46684 087/0115 00-121 PEM AXIS 1 ON
46685 087/0318 00-122 PEM AXIS 2 OFF
46699 088/0118 00-123 OBC CLK ADJ 0.00 TO 20.6
46726 089/0640 00-124 OBC CLK ADJ 20.6 TO 0.00
46726 089/2201 00-125 HRDI MICROLOAD
46740 090/1923 00-126 HRDI MICROLOAD
46740 090/2023 00-127 NBTR NBTR OPS
46743 090/2344 00-128 MLS POWER OFF
46743 091/0024 00-130 HALOE POWER ON
46746 091/0534 00-129 NBTR NBTR OPS
46754 091/1848 00-131 HRDI MICROLOAD
46754 091/1848 00-132 NBTR NBTR OPS
46755 091/2025 00-133 HRDI MICROLOAD
46758 091/2347 00-134 HRDI MICROLOAD
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 31 March 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI PTS PULSE BUS CURRENT MON 46671-4
HRPTSPBIMON Y-HI PTS PULSE BUS CURRENT MON 46686-89
HRPTSPBIMON R-LO PTS PULSE BUS CURRENT MON 46699-702
HRPSEQBIMON Y-HI PTS QUIET BUS CURRENT MON 46702
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None to report this period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 31 March 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46699 088/01:28 (27 Mar 00) from 0.00 to 20.6 msec/day
46726 089/06:47 (29 Mar 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 31 March 2000
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------ -----
Battery performance data not available for this report period.
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
Battery performance data not available for this report period.
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 31 March 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No closure information received
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-005 FOT see list March 00 14170 Generic Late Acq.
SEVNTFW00/13